OpenStudy (anonymous):

How do you do ____ 2+√x+3=x

OpenStudy (anonymous):

Move the 2 to the other side, then square both sides and solve the resulting quadratic.

OpenStudy (anonymous):

\[2+\sqrt{x+3} = x \]\[\implies x -2 = \sqrt{x+3}\]\[\implies(x-2)^2 = x+3\]\[\implies x^2-4x + 4 = x + 3\]\[\implies x^2 - 5x + 1 = 0\] And then use the quadratic formula..